OpenStudy (anonymous):

A boat can travel 27 miles downstream in 1 and 1/2 hours. If the rate of the current is 3mph, find the speed of the boat in still water.

OpenStudy (anonymous):

can you first determine the rate of the boat per hour?

OpenStudy (anonymous):

its 15 mph...

OpenStudy (anonymous):

you mean the speed of the boat in still water?

OpenStudy (anonymous):

yeahhh...

OpenStudy (anonymous):

nvrmnd

OpenStudy (anonymous):

Can you show the steps? How did you get that?

OpenStudy (anonymous):

let speed of the boat in still water= (x)mph speed of boat in downstream =(x+3)mph so distance = speed *time 27=(x+3)*(3/2) 27*2=x*3+3*3 54=3x+9

OpenStudy (anonymous):

54-9=3x 45=3x 15=x

OpenStudy (anonymous):